OpenStudy (reneeshumpert):

What is the solution to the following system of equations? 2x + 3y = 6 x + y = 4 The solution of the system of equations would be? so heres what Im thinking if x+y=4 then those two numbers mutiplied by 2 and 3 would have to equal 6... so theres only two ways to get 4 when using addition. 1+3 and 2+2 ... could somebody help me get on the right track ?

OpenStudy (anonymous):

x= 4 -y 2(4-y) + 3y =6

OpenStudy (reneeshumpert):

explain please

OpenStudy (reneeshumpert):

oh I see what you did

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x=6 y=-4

OpenStudy (reneeshumpert):

8-2y+3y=6 8-5y=6 theeenn?? lol

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yup^

OpenStudy (anonymous):

y=??